Differences

This shows you the differences between two versions of the page.

Link to this comparison view

Both sides previous revision Previous revision
Next revision
Previous revision
operations:agenda2021_09_30 [2021/09/30 13:41]
Simin Salarpour
operations:agenda2021_09_30 [2021/10/14 13:31]
Jayender Kumar
Line 1: Line 1:
-**Chair:**  Simin+**Chair:**  Jay
  
 **Next meeting chair:**  Earl **Next meeting chair:**  Earl
  
-**Attended:**  Jasper, Eloise, JamieJayGuifre+**Attended:**  Japser, Eloise, LimMas, JamieSimin is here
  
-**Apologies:**  Lim+**Apologies:**  Prad
  
 **General Issues:**   **General Issues:**  
  
-  Ho26 is currently pending. Due to structural difficulties, the bearing replacement is being pushed back. Strong winds are also a factor in this process during the spring. December is the most likely month for Ho26 to return to work+Can't send the end expmessage - should be okay now 
-  Hb12 dbbc3 investigation is completed in Bonn. It was delivered there six months ago. It has been fixed and will hopefully be back in Hobart next week. As a result of the changes to some modesthe backend will undergo some adjustments during the next month. Jamie will keep you informed of any changes.+mk5cn timeout errors on yg - at the momentit seems going on. Jamie will look into it. till then, keep an eye on that till now 
 +* delay corrs generally being bad for Hb- nothing much we can do. needs power cyclying more and more and hope that it works. 
 +   
 +   
 +AUM039
  
-**AUM038**+Hobart 12m
  
-Yg: +Notes for wiki after fs-aum command, it is not mentioned in wiki whether you need to select flexbuffhb or flexbuff-hb2, had to confer with Warren (flexbuff-hb2). It is also not mentioned which version of dbbc3hb to run, or whether or not version of dbbchb must also be running. (Eloise - flexbuff-2hb, dbbc3hb should be right choice.
-  * 1805UT Had to sync clocks with fmset because of a marginal error (Pradyumna).+
  
-Hb12: +No sampler “good/bad” readings - sampler summary script (multicast) not installed on pcfshb, and “AUM monitor” on ops8 not working either- the multicast3*.py script is on pcfs-2hbwe can run it from there
-  cannot set fs time without multiple format recorder error from the onsetInvoked 'setupaum' at 1813UT to fix it. Apparently, a benign error? +
-       * It is a chronic problem for both Hb12 and Ke due to their field systems. Using "setupaum" instead of "setupsx" will help in resolving the issue.+
  
-Ke: 
-  * 1730UT - cannot set fs time without multiple format recorder error from the onset. Reset dbbc3, passed the 'setupaum' command, and jive5ab restarted. There was no recording happening. Recording started 1905UT. 
-      * It is a chronic problem for both Hb12 and Ke due to their field systems. Using "setupaum" instead of "setupsx" will help in resolving the issue. 
-  *  2248UT - All samplers had bad stats. Restarted dbbc3ke. Resumed experiment at 2345UT (Pradyumna). 
-  * 0436UT - All samplers gone bad again, I will not reset it for the rest of the experiment. (Lim) 
-       * When samplers are bad, valid data for correlation cannot be collected. It is better to halt the schedule, reset and reconfigure the dbbc, and then continue with the experiment. A clock jump is preferable to bad data. 
  
 +R41019
 +Yarragadee 12m
  
-**AUA080**+A few mk5cn time-out errors which seem benign (Lim). 
 +1156UT mk5cn time-out errors. Back on schedule from 1606UT (Jay). -most prob issue with the module. try to swap the module early on when you see error like this.
  
-Hb12:   
-  * Been running with one bad sampler all day (Warren) 
-       * The issue has been fixed and it is working well now. 
-  * vncviewer from ops8 to pcfshb froze out. I was still trying to diagnose when Jay logged in (Jasper). 
-  * Power outrage at hobart 12 site. can't connect to ipshb either to restart pcfshb. Can't do much remotely. Antenna stowed as of now. Jamie/Warren will have a look in the morning from observatory (Jay). 
  
-Ke: +Si1281 
-  * Been running with one bad sampler all day (Warren)+Hobart 12m
  
 +Bad Luck. Most of Delay Corr went bad as soon as the Exp. started. No point of reconfig. the dbbc again because it will take ~30m and I doubt that it will get good even after that.
 +
 +
 +
 +AOV064
 +Hobart 12m
 +
 +2127UT Delay Corr A D E F gone bad (Lim)
 +Issue is temperature in server room - AC under repair. Temporary ventilation in place but likely to have issues until AC restored. (JMc). First good data after restarting DBBC3 from 0506 (JMc) - one of a time mechanical  issue
 +
 +1313UT -restarted ddbc3 because sampler E was bad. The stats for other samplers worsened after restarting dbbc3. (Prad) - try power cycling it? we only need samplers A,B,C, D. so we can afford sampler E being bad
 +
 +Katherine 12m
 +
 +1655UT “Delay Corr” issue, experiment started late (Lim).
 +Some nuisance errors concerning samplers being marginally below threshold. Issue seems to be transient (JMc) - Jamie will look into log monitor settings for this